Tik-61.181 Bioinformatics Kaski, Mannila, Nikkilä November 8, 2000

Instructions for project report

All written material should be returned to Janne Nikkilä(room B307, janne.nikkila@hut.fi). The material is evaluated and forms part of the grade.

Project work

In order to pass the course you are required to do a relatively small scale project work. Despite of the obligatoriness of the project work, you are encouraged to regard it as an opportunity. You will have a chance to do research in the area of the bioinformatics about any subject you choose and you will get comments and even credits on your work.

So the preferable approach to project work is to sketch the idea of the project work by yourself, write a project plan and when it is accepted proceed to actual work. However, if you have absolutely no idea whatsoever, contact us and we will think about it together.

Project plan

To get feedback on the choice of the project, you are required to write a project plan of about 1/2 A4, describing what you intend to do in your project.

Deadline for project plan: 5.12.2000.

Project Report

Describe your project work in a brief, clear report of about 5 pages. The report should contain

• an abstract,

• a clear introduction that describes the problem setting and goals,

• brief description of the methods and parameters used in the analysis,

• description of the results,

• conclusions, and

• references.

Deadline for written project report: 31.1.2001